OpenStudy (anonymous):

WILL FAN AND MEDAL For a circle of radius 4​ feet, find the arc length subtended by a central angle of 60 degrees. Round to the nearest hundredth.

OpenStudy (welshfella):

arc length = r * O where r = radius and O = the angle in radians so you first need to convert 60 degrees to radians

OpenStudy (anonymous):

The formula for changing degrees to radians is pi/180(deg), right?

OpenStudy (anonymous):

\[\frac{\pi 8\ 60}{360}=4.19 \text{ feet} \]

OpenStudy (anonymous):

pi d is the circumference length. d = 2r = 2*4 = 8